Downpour17:

18q = –11 + 19q What would the come out to if q= 7

Downpour17:

I got 126... I dont know if thats correct or not though

BlasiannPrincess1:

did you subtract?

Downpour17:

I didnt find any subtraction in the problem. The 11 is a negative number

Downpour17:

I really just want to know if my answer is right and if not how do I fix it

BlasiannPrincess1:

if there is a addition sign dat means you subtract but not from the number dat has a variable next to it

Downpour17:

That makes sense

Downpour17:

Kind of

mxddi3:

well, first i did 18(7) and i got 126 as well, but if -11+19(7) has to be equal to 126 then idk how that works.....

BlasiannPrincess1:

Dats not right there is no 7

mxddi3:

it says if q=7

Downpour17:

The variable q is equal to 7

mxddi3:

yeah, 18*7 is 126 but if you replace the variable with 7 on the other side and you have -11+(19*7). 19*7=133 and 133-11=122 so we are left with \[126=122\] but 126 does not equal 122.....

Downpour17:

So did I set up the question wrong?

mxddi3:

you may have. is there a word problem or something you got this from that we can see?

Downpour17:

no I guess the variable. The original question did not give the variable and that was the closest I could come to the answer

BlasiannPrincess1:

Oh den why did you put 18q = -11 +19q?

mxddi3:

what is the original question?

Downpour17:

just 18q = –11 + 19q

mxddi3:

so how did you get 7?

Downpour17:

I used 8 and it went over so I assumed it was 7 because I thought that I was doing the equation wrong

mxddi3:

when you used 8 it went over what? I think i've figured the answer out but I want to see how you are thinking before I tell you my idea

Downpour17:

It went over the variable's quantity. The outcomes are supposed to be equal

mxddi3:

yes. then i have figured it out

Downpour17:

wait is the variable 11?

mxddi3:

ahhh, you got it!

Downpour17:

hey I got it :D

Downpour17:

Thanks for the help!

mxddi3:

do you know how you got that answer?

mxddi3:

and you're welcome

Downpour17:

Yes. I went through numbers from 8 and on until both answers were equal

mxddi3:

ok, good. Let me know if you need help another time!

Downpour17:

Ok thanks!